Find the slope of the line that contains the pair of points...{what does that mean?↙
A(2, 4), B(4, 6)

Answer by Alan3354(69443)

What does what mean?
If you're asking, what does the slope mean:
The slope is the "steepness"
It's the vertical change divided by the horizontal change.
--------------
In the x-y plane, that's diffy/diffx, difference in y over difference in x.
====================================================
A(2,4), B(4,6)
Change in y = 6-2 = 4
Change in x = 4-2 = 2
--------
Slope = 4/2 = 2